The sum of two numbers is 50. If twice the larger is subtracted from four times the smaller, the result is 8. Find the numbers.
Answer by checkley77(12569)

X+Y=50 OR X=50-Y
4X-2Y=8
4(50-Y)-2Y=8
200-4Y-2Y=8
-6Y=8-200
-6Y=-192
Y=-192/-6
Y=32 ANS.
X+32=50
X=50-32
X=18 ANS.
PROOF:
4*18-2*32=8
72-64=8
8=8